A rectangular storage room has dimensions 10 feet by 14 feet by 30 feet.

What is the longest distance a moth could fly in a straight line inside this room? [you’ll need to use a certain Theorem - twice]

Answer by Boreal(15235)

Let length be 30 feet, width 14 feet and height 10 feet
Across the length and width diagonal is sqrt(30^2+14^2)=sqrt(1096)=33.105 feet
-
Across the 3-dimensional triangle, the hypotenuse in the first triangle is now the long side, the short side is the height, and the desired distance is the new hypotenuse.
that is sqrt(1096+10^2)=sqrt(1196)=34.58 feet